OpenStudy (anonymous):

Use Part 1 of the Fundamental Theorem of Calculus to find the derivative of the function. equation in comments

OpenStudy (anonymous):

\[y= \int\limits_{sinx}^{cosx} (1+v^6)^7 dv\]

OpenStudy (anonymous):

chain rule plus ft of c

OpenStudy (anonymous):

replace \(v\) by \(\cos(x)\) then multiply by the derivative of \(\cos(x)\) you get \[-(1+\cos^6(x))^7\sin(x)\]

OpenStudy (anonymous):

then repeat with \(\sin(x)\) but this time subtract

OpenStudy (anonymous):

i get \[-\sin x(1+\cos^6x)^7-\cos x(1+\sin^6x)^7\]

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

yw